OpenStudy (anonymous):

Can you tell which way a parabola will open up just by looking at the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if its y^2=4ax then on right and y^2=-4ax then on left and if its x^2 = 4ay then upwards and if its x^2=-4ay then downwards

OpenStudy (anonymous):

ax^2 + bx + c is also a parabola.

OpenStudy (anonymous):

yeah i know that i am just telling the forms

OpenStudy (anonymous):

IS " if the leading coefficient is greater than zero, the parabola opens upward, and if the leading coefficient is less than zero, the parabola opens downward." a decent answer?

OpenStudy (anonymous):

you can look from y=ax^2+bx+c the state if you want get max or min in the stationary point you have to double differentiation so y''=2a and you know that if y''>0 is min point and parabola is open o if a>0 parabole is open

OpenStudy (anonymous):

in standard form ax^2 + bx + c, if a is positive, opens up if a is negative, opens down

OpenStudy (anonymous):

second derivative? stationary point? opens up if \[a>0\] and down if \[a<0\] oh what mtbender said sorry

OpenStudy (anonymous):

yes, fuzzy...that answer is acceptable. :) let's not nuke it guys ;)

OpenStudy (anonymous):

@fuzzynips IS " if the leading coefficient is greater than zero, the parabola opens upward, and if the leading coefficient is less than zero, the parabola opens downward." a decent answer? Works for me (the sideways one is just a twist).

OpenStudy (anonymous):

y = x^2 is a 'U' y = -x^2 is an 'upside down U'